(p+q)(z-px-qy)=1


1
Expert's answer
2021-01-13T18:28:34-0500

"F=pz-p^2x-pqy+qz-pqx-q^2y-1=0"

"\frac{dx}{F_p}=\frac{dy}{F_q}=\frac{dz}{pF_p+qF_q}=-\frac{dp}{F_x+pF_z}=-\frac{dq}{F_y+qF_z}"

"\frac{dx}{z-2px-qy-qx}=\frac{dy}{-py+z-px-2qy}=\frac{dz}{p(z-2px-qy-qx)+q(-py+z-px-2qy)}="

"=-\frac{dp}{-p^2-pq+p(p+q)}=-\frac{dq}{-pq-q^2+q(p+q)}"


"\frac{dp}{-p^2-pq+p(p+q)}=\frac{dp}{0}\implies p=a=const"

"\frac{dq}{-pq-q^2+q(p+q)}=\frac{dq}{0}\implies q=b=const"

Substitute values into the initial equatio.

Then:

"(a+b)(z-ax-by)=1"

"z-ax-by=c_1"


"\frac{dx}{z-2px-qy-qx}=\frac{dy}{-py+z-px-2qy}=\frac{dp}{0}"

"dy-dx=dp=a"

"y-x=c_2"


The general solution:

"F(c_1,c_2)=F(z-ax-by, y-x)=0"